OpenStudy (anonymous):

HELP PLEASE 29. Find a pair of factors for each number using the difference of two squares. a. 45 b. 77 c. 112

OpenStudy (ajprincess):

Different squares is a^2-b^2 a^2-b^2=(a-b)(a+b) 45 =49-4 =7^2-2^2 =(7-2)(7+2) =5*9 Therefore factors of 45 5 nd 9. 77 =81-4 =9^2-2^2 =(9-2)(9+2) =7*11 Therefore factors of 77 7 nd 11. 112 =121-9 =11^2-3^2 =(11-3)(11+3) =8*14 Therefore factors of 77 8 nd 14.

OpenStudy (ajprincess):

is that clear?

OpenStudy (anonymous):

yes thank you so much. i've been trying to work this all night.

OpenStudy (ajprincess):

u r welcome.